OpenStudy (anonymous):

Last Summer School Question Please HELP WILL FAN AND GIVE MEDAL FOR ANSWER Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions. (3/x)-(3/x+1)-(3/x+2)

OpenStudy (anonymous):

@theEric Can you help me with this one?

OpenStudy (theeric):

\[\left(\frac{3}{x}\right)-\left(\frac{3}{x}+1\right)-\left(\frac{3}{x}+2\right)\] Any idea where to start?

OpenStudy (anonymous):

oh the 1 and 2 are with the x's under the bar..

OpenStudy (theeric):

(3/x)-(3/(x+1))-(3/(x+2))? Okay! :)

OpenStudy (anonymous):

Yup! So I know they all need the same denominator

OpenStudy (theeric):

\[\left(\frac{3}{x}\right)-\left(\frac{3}{x+1}\right)-\left(\frac{3}{x+2}\right)\]

OpenStudy (anonymous):

Yeah that's right so you would.... multiply the first one by (x+1) and (x+2)?

OpenStudy (theeric):

Yeah, that's the basic idea... I looked on the internet quick to try to find a shortcut, but found none so far. We can do it that way, but it'll be messy! And we always multiply by \(1\), so we don't change the actual amount! (Anything times \(1\) is itself.) And\[1=\frac{x}{x}=\frac{x+1}{x+1}=\frac{x+2}{x+2}\]:)

OpenStudy (anonymous):

whoa okay... uh what next then?

OpenStudy (theeric):

So you know we're headed towards a common denominator of \((x)(x+1)(x+2)\), so you want to write that down. \[\left(\frac{3}{x}\right)\left(\frac{x+1}{x+1}\right)\left(\frac{x+2}{x+2}\right)-\left(\frac{3}{x+1}\right)-\left(\frac{3}{x+2}\right)\]That's what you said! Now

OpenStudy (anonymous):

Okay I got \[\left( 3x ^{2}+9x+6 \div x(x+1)(x+2)\right)\]

OpenStudy (theeric):

Let me see what I get.

OpenStudy (theeric):

Hmm... We got a different answer, but I don't know which of us did the wrong thing. So! Did you get\[\frac{3(x+1)(x+2)-3(x)(x+2)-3(x)(x+1)}{x(x+1)(x+2)}\] along the way?

OpenStudy (anonymous):

After going back and doing that to all of them yes I did. for my final answer I got...\[-3x ^{2}/x(x+1)(x+2)\]

OpenStudy (theeric):

Alright! Now there's just one slight difference. I would like you to look back at one part in your work, which will be enlarged here:\[\small\frac{3\Large (x+1)(x+2)\small-3(x)(x+2)-3(x)(x+1)}{x(x+1)(x+2)}\] Do you find that \((x+1)(x+2)=x^2+3x+2\)?

OpenStudy (anonymous):

Yes I did!

OpenStudy (theeric):

Alright! :) How did you cancel that \(2\) out?

OpenStudy (theeric):

Because I didn't, in my work.

OpenStudy (anonymous):

Oh what I did was multiply them all out... And I'm mistaken. I canceled it out after multiplying them all out and cancelling out the 6 with at 6x... Whoops

OpenStudy (theeric):

Haha, easy mistake! So, how dos that affect your final answer? I did cancel out all my terms with a plain \(x\) in them. I got \[\frac{3(2-x^2)}{x(x+1)(x+2)}=\frac{6-6x^2}{x(x+1)(x+2)}\]

OpenStudy (anonymous):

Yes that is what I got after going back through!

OpenStudy (theeric):

Cool! Now, about the denominator. Should we FOIL that?

OpenStudy (theeric):

I don't know if that is making it simpler or not.

OpenStudy (anonymous):

Uhhh... Yes? Well I guess what I would do would be to foil out (x+1)(x+2) and then multiply x with the answer to that?

OpenStudy (theeric):

Okay!

OpenStudy (anonymous):

So it would be \[3x ^{3}+3x ^{2}+3x\]

OpenStudy (anonymous):

?

OpenStudy (theeric):

\((x+1)(x+2)=(x\times x)+(x\times 2)+(1\times x)+(1\times 2)\\=x^2+2x+x+2\\=x^2+3x+2\)

OpenStudy (theeric):

\((x)(x^2+3x+2)=(x\times x^2)+(x\times 3x)+(x\times 2)\\=x^3+3x^2+2x\)

OpenStudy (theeric):

Did I do it right? Or can you find your mistake?

OpenStudy (anonymous):

Oh! Rookie mistake I meant 2!

OpenStudy (anonymous):

2*1=2 wow can you tell I'm tired? haha

OpenStudy (theeric):

Haha, tiny mistakes...

OpenStudy (theeric):

Oh gosh.. You did say this is your last question, though! Until you get more familiar and relaxed with algebra, it can be exhausting in itself.

OpenStudy (theeric):

on it's own, I mean.

OpenStudy (anonymous):

Yes one of them! I think there are two more after this one but they shouldn't be that big of a deal... Or trouble... If you know what I mean

OpenStudy (theeric):

So, now you need to look through your work and make sure you know how you got from each step to the next (and which steps you should erase because they had mistakes).

OpenStudy (theeric):

I know what you mean. OpenStudy will be here. And you seem to know what you're doing anyway.

OpenStudy (theeric):

By the way, before you simplified, the denominator was \(x(x+1)(x+2)\). That, the denominator, cannot be \(0\) or else the expression is undefined. I think that is what the problem wants to know when it asks for "restrictions".

OpenStudy (theeric):

So, for the restrictions, you have to say what \(x\) cannot be.

OpenStudy (anonymous):

Sorry! my wifi went haywire! And I would say that x cannot equal -1 or -2 because then you will be dividing by 0.

OpenStudy (theeric):

Ah, that stinks! Right! There's one more thing \(x\) can be!

OpenStudy (anonymous):

x can't be 0...?

OpenStudy (theeric):

Right. :)

OpenStudy (anonymous):

Okay! Thank you so much!

OpenStudy (theeric):

You're welcome! Good luck with the sentences!

OpenStudy (anonymous):

Thank you! Do you have time for one more question?

OpenStudy (theeric):

Sure! Post it! Then send me a link to or it use the "@" thing.